OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2.

OpenStudy (anonymous):

Do you know how to get the equation given those two values? ^

OpenStudy (anonymous):

no

OpenStudy (anonymous):

http://www.algebra.com/algebra/homework/Quadratic-relations-and-conic-sections.faq.question.83012.html Read through this. O.o it might help.

hero (hero):

Insert points (0 , 2) and (x, -2) into the formula \((x - x_1)^2 + (y - y_1)^2 = (x - x_2) + (y - y_2)^2\) to get: \((x - 0)^2 + (y - 2)^2 = (x - x) + (y - (-2))^2\) \(x^2 + (y - 2)^2 = 0 + (y + 2)^2\) \(x^2 + y^2 - 4y + 4 = y^2 + 4y + 4\) Finish isolating y from there.

OpenStudy (anonymous):

y^2=2x

hero (hero):

\[x^2 + \cancel{y^2} - 4y + \cancel{4} = \cancel{y^2} + 4y + \cancel{4}\]

OpenStudy (anonymous):

Okay that's not an option though

hero (hero):

You're not done simplifying it either.

hero (hero):

What did you get for the final simplification?

hero (hero):

The next step would be to add 4y to both sides to get: \(x^2 = 4y + 4y\)

hero (hero):

Obviously, \(4y + 4y = 8y\) so \(x^2 = 8y\)

OpenStudy (anonymous):

y^2=8x

hero (hero):

Based on the given information, \(x^2 = 8y\) is the correct result.

hero (hero):

I apologize if it isn't one of your options.

OpenStudy (anonymous):

Yes I see but it's not an option, here are the options: y^2 = 2x y = 1/2x^2 y^2 = 8x y = 1/8x^2

hero (hero):

Ah, but it is an option.

OpenStudy (anonymous):

y =1/8x^2?

hero (hero):

Yes, exactly.

OpenStudy (anonymous):

Cool (((:

hero (hero):

\(x^2 = 8y\) Divide both sides by 8 \(\dfrac{x^2}{8} = y\) \(y = \dfrac{x^2}{8}\) \(y = \dfrac{1}{8}x^2\)

hero (hero):

\(\dfrac{x^2}{8}\) means the same as \(\dfrac{1}{8}x^2\). Try to understand why

OpenStudy (anonymous):

Ahh got it (: Thank you!!